OpenStudy (anonymous):

A bricklayer and an electrician together spend 90 hours working on a new house. If the bricklayer earns $12 per hour, the electrician earns $16 an hour, and the owner pays them a total of $1350 for their work, how many hours does each worker spend on the house?

OpenStudy (saifoo.khan):

Let x=amount of time spent by the bricklayer And let y=amount of time spent by the electrician Now we are told the following x+y=90----------------------------eq1 and 12x+16y=1350 ($ understood)-----------------------eq2 Multiply eq1 by 12: 12x+12y=1080--------------------eq1a Subtract eq1a from eq2: 4y=270 divide each side by 4 y=67.5 hours----------------------------amount of time spent by the electrician Substitute y=67.5 into eq1: x+67.5=90 add 67.5 to each side x+67.5-67.5=90-67.5 collect like terms x=22.5 hours-------------------------------time spent by the bricklayer

OpenStudy (anonymous):

thank you